OpenStudy (anitas.):

Two lines, A and B, are represented by the equations given below: Line A: x + y = 2 Line B: 2x + y = 4 Which statement is true about the solution to the set of equations? There are infinitely many solutions. There are two solutions. There is one solution. There is no solution.

OpenStudy (anitas.):

@Mehek14 plz show me how to do this???

OpenStudy (michele_laino):

hint: please solve the first equation for \(y\)

OpenStudy (anitas.):

? how do i solve it?

OpenStudy (michele_laino):

if I add \(-x\) to both sides, I get: \(x-x+y=2-x\) please simplify

OpenStudy (michele_laino):

what is \(x-x=...?\)

OpenStudy (anitas.):

y

OpenStudy (michele_laino):

I think that: \(x-x=0\), right?

OpenStudy (anitas.):

oh yes i gues so

OpenStudy (michele_laino):

so, we can write: \(y=2-x\) am I right?

OpenStudy (michele_laino):



OpenStudy (anitas.):

so there is one solution right?

OpenStudy (michele_laino):

please wait next: if I subtract \(2x\) from both sides of the \(second\) equation, I get: \[\Large 2x - 2x + y = 4 - 2x\] please simplify

OpenStudy (michele_laino):

what is \(2x-2x=...?\)

OpenStudy (anitas.):

2x=4−y?

OpenStudy (michele_laino):

correct! :) nevertheless such equation is not useful. An equation which can be useful is: \[\Large y = 4 - 2x\]

OpenStudy (michele_laino):



OpenStudy (michele_laino):

now, from first equation we got \(y=2-x\) whereas from second equation we got \(y=4-2x\), then we can write this: \(\Large 2-x=4-2x\)

OpenStudy (michele_laino):

please solve that equation with respect to \(x\)

OpenStudy (michele_laino):

if we draw both of those lines, using the coordinates \((x,y)\), we get:

OpenStudy (michele_laino):

as we can see, we have only one intersection point, so what can you conclude?

OpenStudy (anitas.):

that we got one solution

OpenStudy (michele_laino):

that's right!

OpenStudy (anitas.):

thank you very much for explaining this whole thing to me @Michele_Laino it's very nice of you!!! :D

OpenStudy (michele_laino):

thanks!! :D

OpenStudy (anonymous):

subtract first equation from second equation and find x and then corresponding y.